Circular No. 8018

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hh IN NGC 6946 C. J. Stockdale, Naval Research Laboratory (NRL) and National Re-

search Council; R. A. Sramek and M. Rupen, National Radio Astronomy Observatory (NRAO); K. W. Weiler, NRL; S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology; N. Panagia, European Space Agency and Space Telescope Science Institute; D. Pooley and W. Lewin, Massachusetts Institute of Technology; and S. Myers and G. Taylor, NRAO, report the detection of radio emission from the type-II supernova 2002hh (IAUC 8005, 8007) with the Very Large Array in ‘C’ configuration: “Radio-flux densities on Nov. 15.25 UT of 0.60 ± 0.10 mJy at 22.485 GHz (wavelength 1.3 cm) and 0.81 ± 0.08 mJy at 8.435 GHz (wavelength 3.6 cm) were observed at α = 20^{h}34^{m}44^{s}.25, δ = +60^{o}07′19′′.4 (equinox 2000.0; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (end figures 44^{s}.29, 19′′.0; IAUC 8005). On Nov. 8.0, no radio emission was detected at that position, with 3σ upper limits of 0.44 mJy at 22.485 GHz and 0.19 mJy at 8.435. Although few conclusions can be drawn from such sparse data, the apparently optically thin character of the radio emission, if confirmed, may indicate that the circumstellar interaction is weak and is evolving unusually rapidly, as was seen for, e.g., SN 1981K. Radio observations are continuing.”

LEONID METEORS 2002 Visual reports indicate peaks in Leonid activity around Nov. 19.174

(primary peak) and 19.45 UT (secondary), agreeing with predictions by P. Jenniskens (2001, WGN 29, 165) for Nov. 19.172 (for material from the return of comet 55P in 1767) and 19.426 (1866 return), and by McNaught and Asher (2002, WGN 30, 132) for Nov. 19.164 and 19.440. Visual rates of up to 10–25 Leonids/min around Nov. 19.14–19.18 have been received from Á. Kereszturi, K. Sárneczky, I. Tepliczky, and colleagues in Hungary; A. Baransky at Bucha, Ukraine; J. D. Shanklin at St. Maximin, Provence, France; and K. Hornoch at Lelekovice, Czech Republic. A. Hale, Cloudcroft, NM, reports a peak rate of ≈ 150/hr during Nov. 19.438–19.458.

CENTRAL BUREAU COMPUTER SERVICE Subscribers who use the Central Bureau’s Computer Service are ad-

vised that the host names for the computers were changed to cfaps1.cfa.

harvard.edu and cfaps8.cfa.harvard.edu (IP addresses are unchanged).

2002 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green